OpenStudy (anonymous):

The cylinder shown below has a radius of 9 inches and a height of 31 inches.

OpenStudy (anonymous):

What is the approximate lateral area of the cylinder?

OpenStudy (anonymous):

279

OpenStudy (anonymous):

My choices are 876.86 in² 1,131.43 in² 1,753.71 in² 2,262.86 in² It also says use 22/7 for pi.

OpenStudy (anonymous):

the formula for finding the lateral area of a cylinder is L.A.= 2*pi*radius*height so lets fill that in L.A. = 2*22/7*9*31 L.A. = 1753.71 in² if you want to know how you get to that formula, try to imagine cutting open the cylinder, so it forms a rectangle. in that case you need to know the width and height of the rectangle: we know that the height is 31in the width is the same as the perimeter of the circle (top of the cylinder) to find the perimeter of a circle: 2*pi*radius so you multiply those, which gives you: L.A. = 2*pi*radius*height